# Compelling ReLU Networks to Exhibit Exponentially Many Linear Regions at Initialization and During Training

### v6

@@ Abstract
In a neural network with ReLU activations, the number of piecewise linear regions in the output can grow exponentially with depth. However, this is highly unlikely to happen when the initial parameters are sampled randomly, which therefore often leads to the use of networks that are unnecessarily large. To address this problem, we introduce a novel parameterization of the network that restricts its weights so that a depth $d$ network produces exactly $2^{d}$ linear regions at initialization and maintains those regions throughout training under the parameterization. This approach allows us to learn approximations of convex, one-dimensional functions that are several orders of magnitude more accurate than their randomly initialized counterparts. We further demonstrate how to extend our approach to multidimensional and non-convex functions, allowing it to replace the dense layers in other networks; preliminary improvements are shown for image classification on CIFAR-10 and ImageNet.
Machine Learning, ICML
@@ 1 Introduction
Beyond complementary advances in areas like hardware, storage, and networking, the success of neural networks is primarily due to their ability to efficiently capture and represent nonlinear functions (Gibou et al., 2019). In a neural network, the goal of an activation function is to introduce nonlinearity between the network’s layers so that the network does not simplify to a single linear function. The rectified linear unit (ReLU) has a unique interpretation in this regard. Since it either deactivates a neuron or acts as an identity, the resulting transformation on each individual input remains linear. However, each possible configuration of active and inactive neurons can produce a unique linear transformation over a particular region of input space. The number of these activation patterns and their corresponding linear regions provides a way to measure the expressivity of a ReLU network [^1] See Appendix A.2 for definitions of terms like linear regions, activation patterns, and activation regions. and can theoretically scale exponentially with the depth of the network (Montufar et al., 2014a; Serra et al., 2018). Hence deep architectures may outperform shallow ones.
Surprisingly, though, a sophisticated theory of how to best encode functions into ReLU networks is lacking, and in practice, adding depth is often observed to help less than one might expect from this exponential intuition. Lacking more advanced theory, practitioners typically use random parameter initialization and gradient descent, the drawbacks of which often lead to extremely inefficient solutions. Indeed, Hanin & Rolnick (2019) show a rather disappointing bound for randomly initialized networks: the average number of linear regions formed at initialization is invariant with depth. Whether using a deep or shallow architecture, only the total number of neurons matters. They further observed that gradient descent has a difficult time creating new activation regions so that their bounds approximately held after training. As we will discuss later, the total number of linear regions is not a “local” property in parameter space that gradient descent can directly optimize (see 5). Gradient descent is also prone to redundancy; Frankle & Carbin (2019) show how around 95% of weights may ultimately be eliminated from a network without significantly degrading accuracy.
@@
  • Extensions to non-convex and multidimensional functions, and use of our method as a drop-in replacement for dense layers in arbitrary networks (Section 5; Appendices A.10A.11)
While an exponential increase in the expressiveness of a ReLU network does not necessarily imply an exponential increase in performance, one may intuitively expect a substantial benefit, and our results bear this out, yielding error values orders of magnitude lower than a traditionally-trained network of equal size for simple test problems (Section 4), and quickly outpacing such networks’ performance on more practical test cases such as ImageNet (Appendix A.11). We emphasize that while the preliminary mathematical analysis in this work pertains to a specific ReLU construction, this construction can be repeated and combined in useful ways to obtain performance improvements on larger networks (Section 5). We conclude in Section 6 with discussions on extending the mathematical framework of our method to arbitrary ReLU networks and other types of networks, which would have significant practical utility.
@@ 2.1 Function Approximation
The same effect can be achieved with non-symmetric triangle waves (Huchette et al., 2023) (or any shape that sufficiently “folds” the input space (Montufar et al., 2014b)). Our reparameterization strategy (Section 3) focuses on non-symmetric triangle waves.
@@
The $x^{2}$ approximation is used by other theoretical works as a building block to guarantee exponential convergence rates in more complex systems. Perekrestenko et al. (2018) construct a multiplication gate via the identity $(x+y)^{2}=x^{2}+y^{2}+2xy$. The squared terms can all be moved to one side, expressing the product $xy$ as a linear combination of squared terms. They then further assemble these multiplication gates into an interpolating polynomial, which can have an exponentially decreasing error when the interpolation points are chosen to be the Chebyshev nodes. Polynomial interpolation does not scale well into high dimensions, so this and papers with similar approaches will usually come with restrictions that limit function complexity: Wang et al. (2018) requires low input dimension, Montanelli et al. (2020) uses band limiting, and Chen et al. (2019) approximates low-dimensional manifolds. These works all make use of a fixed representation of $x^{2}$. If our networks were substituted in for the $x^{2}$ approximation, these works would provide theoretical guarantees about the capabilities of the resulting model. Even though the approximation rates will not scale well with input dimension, they serve as a bound that can be improved upon. In Section 5 we further elaborate on how to use our networks to represent higher-dimensional or non-convex functions.
@@ 2.2 Neural Network Initialization
Our work seeks to improve network initialization by making use of explicit theoretical constructs. This stands in sharp contrast the current standard approach, which treats neurons homogeneously. Two popular initialization methods implemented in PyTorch are the Kaiming (He et al., 2015) and Xavier initialization (Glorot & Bengio, 2010). They use weight values that are randomly sampled from distributions defined by the input and output dimension of each layer. Aside from sub-optimal approximation power associated with random weights, a common issue is that the initial weights and biases in a ReLU network can cause every neuron in a particular layer to output a negative value. The ReLU activation then sets the output of that layer to 0, blocking any gradient updates. This is referred to as the dying ReLU phenomenon (Qi et al., 2024; Nag et al., 2023). Worryingly, as depth goes to infinity, the dying ReLU phenomenon becomes increasingly likely (Lu et al., 2020). Several papers propose solutions: Shin & Karniadakis (2020) use a data-dependent initialization, while Singh & Sreejith (2021) introduce an alternate weight distribution called RAAI that can reduce the likelihood of the issue and increase training speed. We observed during our experiments that RAAI greatly reduces, but does not eliminate, the likelihood of dying ReLU. Our approach enforces a specific network structure that does not collapse in this manner.
@@ 3 Initialization and Pretraining Construction
We begin this section by describing our ReLU network reparameterization strategy that ensures an exponential number of linear regions with respect to depth. We focus on architecting the weights of a 4-neuron-wide, depth $d$ ReLU network; while this is a constraint of the present manuscript, we will show that despite this, we can effectively apply our construction to a wide variety of examples.
@@
where $10<a_{i}<$ 1. This produces a triangular shape with a peak at $x=a_{i}$ and with both endpoints satisfying $y=0$. Each layer considered in isolation would compute these if directly fed the input signal. $T_{i}(x)$’s derivatives are the piecewise constant functions:
$T^{\prime}_{i}(x)=\begin{cases}\frac{1}{a_{i}}&0<x<a_{i}\\ \frac{1}{1-a_{i}}&a_{i}<x<1\end{cases}$ (1)
In a deep network, the layers feed into each other, composing their respective triangle functions:
$W_{i}(x)=\bigcirc_{j=0}^{i}T_{j}(x)=T_{i}(T_{i-1}(...T_{0}(x)))$ (2)
These triangle waves will have $2^{i}$ linear regions, doubling with each layer. The output of the network in pretraining will be a weighted sum over the triangle waves formed at each layer. Assuming the network to be infinitely deep, we have
$F(x)=\sum_{i=0}^{\infty}s_{i}W_{i}(x)$ (3)
where $s_{i}$ are scaling coefficients on each of the composed triangular waveforms $W_{i}$.
@@
To encode these functions into the weights of a ReLU network, we begin with triangle functions (see the right subnetwork in Figure 2). Its maximum output is $1$ at the peak location $a\in(0,1)$. Neuron $t_{1}$ simply preserves the input signal. Meanwhile, $t_{2}$ is negatively biased, deactivating it for inputs less than $a$. Subtracting $t_{2}$ from $t_{1}$ changes the slope at the point where $t_{2}$ begins outputting a nonzero signal. The weight $-1/(a-a^{2})=-(1/a+1/(1-a))$ is picked to completely negate $t_{1}$’s positive influence, and then produce a negative slope. When these components are combined into a deep network, the individual triangles they form will be composed with those earlier in the network, but the two neurons will still behave analogously. Considering the output waveform of $t_{1}$ and $t_{2}$ neurons over the entire input domain $[0,1]$ and at arbitrary depth, we have that $t_{1}$ neurons are always active, outputting complete triangle waves. In contrast, $t_{2}$ neurons are deactivated for small inputs, so they output an alternating sequence of triangles and inactive regions (see Figure 3).
@@
Naively using the output of one triangle generator (from Figure 2) as input to the next would form a $1\rightarrow 2\rightarrow 1\rightarrow 2\rightarrow 1$ shape, but this is unnecessarily deep and irregular. We can replicate the one-dimensional function composition in the hidden layers of our network (see Figure 2 top right). Any outgoing weight from $t_{1}$ or $t_{2}$ is shared; every neuron taking in a triangle wave as input does so by combining $t_{1}$ and $t_{2}$ in the same proportion. In this way, we can avoid having to use extra intermediate neurons. This is evident in Figure 3 where the waves output by $t_{1}$ and $t_{2}$ neurons possess identical slopes in regions where both neurons are active.
There are two other neurons in the pretraining network’s layers, each with a specific role. The accumulation neuron (marked as “sum” in Figure 2) maintains a weighted sum of all previous triangle waves through each layer. This is similar to a residual connection (He et al., 2016), except that no other neurons take the sum as input (its output is convex, and therefore not rapidly oscillating, and detrimental to making more linear regions). If the sum neuron were naively implemented, it would multiply the $t_{1}$ and $t_{2}$ weights by the sum coefficients. Based on Theorem 4, these coefficients will be exponentially decaying, so learning these weights directly may cause conditioning issues. Instead, all weights in each layer are multiplied by the ratios between successive scaling coefficients, so that the outputs of $t_{1}$ and $t_{2}$ neurons decay in amplitude in each layer. A conventional bias will have no connections to prior layers, so it will be unable to adjust to the rescaling neuron outputs without having to learn exponentially small values. Therefore, a fourth neuron is configured to output a constant signal, so the weight other neurons place on this output can replace their bias. The bias neuron will connect to the previous layer’s bias neuron so that the constant signal can scale down gradually with each layer. We handle this scaling explicitly, rather than relying on regularizers like batch norm (Ioffe & Szegedy, 2015) or layer norm (Lei Ba et al., 2016), partially to preserve the mathematical transparency of our model, and partially because this solution arises naturally from our derivations.
@@ 3.2 Differentiable Model Output
Given the rapid oscillations of the triangle waves formed at each layer, the network will output a fractal with many choices of scaling parameters. This would be poorly predictive of unseen data points generated from a smooth curve. Our main mathematical result addresses this issue by forcing differentiability of the network output. We do so by making the sum of the derivatives of the triangle waves converge. For this class of functions, the peak locations of the triangle waves uniquely determine the scales with which to sum them. Mathematical analysis and sufficient conditions for differentiability in the limit are in Appendix A.3.
@@
$s_{i+1}=s_{i}(1-a_{i+1})a_{i+2}$ (4)
@@ 4 One-Dimensional Experiments
Since our network construction applies directly to one dimensional convex functions, this will be the focus of this section (we cover more difficult functions in Sections 56 and A.10A.11) The aim of these experiments is twofold: (1) we would like to determine how to learn the most effective function representations possible, and (2) to explore how the utilization of an increased number of linear regions can affect a network’s ability to capture underlying nonlinearity in its training data. To demonstrate that our networks can learn function representations that better utilize depth, we benchmark against PyTorch’s (Paszke et al., 2019) default settings (nn.linear() uses Kaiming initialization), as well as the RAAI distribution from Singh & Sreejith (2021), and produce errors that are orders of magnitude lower than both. We also train several permutations of our reparameterized networks, which validate that our pretraining and our differentibility constraints indeed facilitate smoother navigation of the loss landscape. Lastly, we conduct a second round of tests to determine if pretrained networks display an enhanced predictive capacity on unseen data points, as might be expected if they can leverage greater nonlinearity in their outputs.
@@ 4.1 Experimental Setup
All models are trained using Adam (Kingma & Ba, 2017) as the optimizer with a learning rate of 0.001 for 1,000 epochs to ensure convergence (see Appendix A.9 for a brief analysis of different learning rates). Each network is four neurons wide with five hidden layers, along with a one-dimensional input and output. The loss function used is the mean squared error, and the average and minimum loss are recorded for 30 models of each type. The networks using triangle based parameterizations share a common set of starting locations where the triangle peaks and scaling parameters are chosen according to our main theorem (Equation 4). In our tables, “pretraining skipped” is a network that initializes with our method, but only trains in the standard parameterization. “differentiability not enforced” is a network that initializes with our parameterization, but then trains the scales of its triangle waves as an extra degree of freedom (so that the final output is not necessarily differentiable). “differentiability enforced” corresponds to the main variant of our method as we have presented it thus far. The four curves we approximate are $x^{3}$, $x^{11}$, $\tanh(3x)$, and a quarter period of a sine wave. To approximate the sine and the hyperbolic tangent, the triangle waves are added to the line $y=x$. For the other approximations, the waves are subtracted. This requires the first scaling factor to be $a_{0}*a_{1}$ instead of $(1-a_{0})*a_{1}$. The first set of data is 500 evenly spaced points on the interval $[0,1]$ for each of the curves. This is chosen to be very dense deliberately, to try to evoke the most accurate representations from each network. We determine from these tests that pretraining with differentiability enforced produces the best results, so we compare it to standard networks in our second set of experiments. We use a second set of data consisting of only 10 points, with a test set of 10 points spaced in between so as to be as far away from learned data as possible. This set of experiments evaluates the predictive capacities of the networks on unseen data.

### v7

@@ Abstract
In a neural network with ReLU activations, the number of piecewise linear regions in the output can grow exponentially with depth. However, this is highly unlikely to happen when the initial parameters are sampled randomly, which therefore often leads to the use of networks that are unnecessarily large. To address this problem, we introduce a novel parameterization of the network that restricts its weights so that a depth $d$ network produces exactly $2^{d}$ linear regions at initialization and maintains those regions throughout training under the parameterization. This approach allows us to learn approximations of convex, one-dimensional functions that are several orders of magnitude more accurate than their randomly initialized counterparts. We further demonstrate a preliminary extension of our construction to multidimensional and non-convex functions, allowing the technique to replace traditional dense layers in various architectures.
ReLU networks, network initialization, linear regions, activation regions
[…]
mytheoremm
[…]
@@ 1 Introduction
Beyond complementary advances in areas like hardware, storage, and networking, the success of neural networks is primarily due to their ability to efficiently capture and represent nonlinear functions (Gibou et al., 2019). In a neural network, the goal of an activation function is to introduce nonlinearity between the network’s layers so that the network does not simplify to a single linear function. The rectified linear unit (ReLU) has a unique interpretation in this regard. Since it either deactivates a neuron or acts as an identity, the resulting transformation on each individual input remains linear. However, each possible configuration of active and inactive neurons can produce a unique linear transformation over a particular region of input space. The number of these activation patterns and their corresponding linear regions provides a way to measure the expressivity of a ReLU network [^1] See Appendix A.2 for definitions of terms like linear regions, activation patterns, and activation regions. and can theoretically scale exponentially with the depth of the network (Montufar et al., 2014; Serra et al., 2018). Hence deep architectures may outperform shallow ones.
Surprisingly, though, a sophisticated theory of how to best encode functions into ReLU networks is lacking, and in practice, adding depth is often observed to help less than one might expect from this exponential intuition. Lacking more advanced theory, practitioners typically use random parameter initialization and gradient descent, the drawbacks of which often lead to extremely inefficient solutions. Indeed, Hanin & Rolnick (2019) show a rather disappointing bound for randomly initialized networks: the average number of linear regions formed at initialization is invariant with depth. Whether using a deep or shallow architecture, only the total number of neurons matters. They further observed that gradient descent has a difficult time creating new activation regions, and thus their bounds approximately held after training. As we will discuss later, the total number of linear regions is not a “local” property in parameter space that gradient descent can directly optimize (see Figure 5). Gradient descent is also prone to redundancy; for instance, Frankle & Carbin (2019) show how around 95% of weights may ultimately be eliminated from a network without significantly degrading accuracy.
@@
  • Extensions to non-convex and multidimensional functions, and use of our method as a drop-in replacement for dense layers in arbitrary networks (Sections 5 and 6; Appendices B.2B.3)
While an exponential increase in the expressiveness of a ReLU network does not necessarily imply an exponential increase in performance, one may intuitively expect a substantial benefit, and our results bear this out, yielding error values orders of magnitude lower than a traditionally-trained network of equal size for simple test problems (Section 4), and outpacing such networks’ performance during pretraining on more practical test cases such as ImageNet (Section 6). We emphasize that while the preliminary mathematical analysis in this work pertains to a specific ReLU construction, this construction can be repeated and combined in useful ways to obtain performance improvements on larger networks (Section 5). We conclude in Section 7 with discussions on extending the mathematical framework of our method to arbitrary ReLU networks and other types of networks, which would have significant practical utility.
@@ 2.1 Function Approximation
The same effect can be achieved with non-symmetric triangle waves (Huchette et al., 2023) (or any shape that sufficiently “folds” the input space (Montufar et al., 2014)). Our reparameterization strategy (Section 3) focuses on non-symmetric triangle waves.
@@
The $x^{2}$ approximation is used by other theoretical works as a building block to guarantee exponential convergence rates in more complex systems. Perekrestenko et al. (2018) construct a multiplication gate via the identity $(x+y)^{2}=x^{2}+y^{2}+2xy$. The squared terms can all be moved to one side, expressing the product $xy$ as a linear combination of squared terms. They then further assemble these multiplication gates into an interpolating polynomial, which can have an exponentially decreasing error when the interpolation points are chosen to be the Chebyshev nodes. Polynomial interpolation does not scale well into high dimensions, so this and papers with similar approaches will usually come with restrictions that limit function complexity: Wang et al. (2018) requires low input dimension, Montanelli et al. (2020) uses band limiting, and Chen et al. (2019) approximates low-dimensional manifolds. These works all make use of a fixed representation of $x^{2}$. If our networks were substituted in for the $x^{2}$ approximation, these works would provide theoretical guarantees about the capabilities of the resulting model. Even though the approximation rates will not scale well with input dimension, they serve as a bound that can be improved upon. In Section 5, we further elaborate on how to use our networks to represent higher-dimensional or non-convex functions.
@@ 2.2 Neural Network Initialization
Our work seeks to improve network initialization by making use of explicit theoretical constructs. This stands in sharp contrast current standard approaches, which treat neurons homogeneously. Two popular initialization methods implemented in PyTorch are the Kaiming (He et al., 2015) and Xavier initialization (Glorot & Bengio, 2010). They use weight values that are randomly sampled from distributions defined by the input and output dimension of each layer. Aside from sub-optimal approximation power associated with random weights, a common issue is that the initial weights and biases in a ReLU network can cause every neuron in a particular layer to output a negative value. The ReLU activation then sets the output of that layer to 0, blocking any gradient updates. This is referred to as the dying ReLU phenomenon (Qi et al., 2024; Nag et al., 2023). Worryingly, as depth goes to infinity, the dying ReLU phenomenon becomes increasingly likely (Lu et al., 2020). Several papers propose solutions: Shin & Karniadakis (2020) use a data-dependent initialization, while Singh & Sreejith (2021) introduce an alternate weight distribution called RAAI that can reduce the likelihood of the issue and increase training speed. We observed during our experiments that RAAI greatly reduces, but does not eliminate, the likelihood of dying ReLU. Our approach enforces a specific network structure that does not collapse in this manner.
@@ 3 Initialization and Pretraining Construction
In this section we describe our ReLU network reparameterization strategy that ensures an exponential number of linear regions with respect to depth. We focus on architecting the weights of a 4-neuron-wide, arbitrary depth ReLU network; while this is a constraint of the present manuscript, we will show that despite this, we can effectively apply our construction to a wide variety of examples.
@@
$T_{i}(x)=\begin{cases}\frac{x}{a_{i}}&0\leq x\leq a_{i}\\ 1-\frac{x-a_{i}}{1-a_{i}}&a_{i}\leq x\leq 1\end{cases},$
where $10<a_{i}<$ 1. This produces a triangular shape with a peak at $x=a_{i}$ and zeros at $x=0$ and $x=1$. These functions are implemented by two ReLU neurons in each layer. In a deep network, each layer composes its triangle function with the output of the layer before it. The result is that each layer computes the following triangular waveform:
$W_{i}(x)=\bigcirc_{j=0}^{i}T_{j}(x)=T_{i}(T_{i-1}(...T_{0}(x))).$ (1)
These triangle waves will have $2^{i}$ linear regions, doubling with each layer. By allocating an extra neuron in each layer, our construction will take a weighted sum over each layer’s triangle wave, using them as a deep basis to build the output. In the infinite depth limit, the network output takes the form:
$F(x)=\sum_{i=0}^{\infty}s_{i}W_{i}(x),$ (2)
where $s_{i}$ are scaling coefficients on each of the composed triangular waveforms $W_{i}$. $F(x)$ would be a parabola if we exactly followed the related work. $F(x)$ could also end up being a fractal curve if the $s_{i}$ are chosen to be too large. Choosing $s_{i}$ carefully to regularize $F(x)$ is a major part of our construction.
@@
Now we will discuss how to encode these functions into the weights of a ReLU network, beginning with a network for a single triangle function (see the upper left diagram in Figure 2). Each triangle function has two nonzero linear pieces, requiring two ReLU neurons. For simplicity, we will set both neurons’ input weights to $1$. We will arrange for neuron $t_{1}$ to be active for all inputs, and since the maximum output should be $1$ at the peak location $a\in(0,1)$, neuron $t_{1}$ will have an outgoing weight of $1/a$. $x=a$ is where the slope should change, so neuron $t_{2}$ will be biased by $-a$ to suppress it for $x<a$. Since both neurons are active for $x>a$, neuron $t_{2}$ will need to be weighted by $-(1/a+1/(1-a))=-1/(a-a^{2})$; this weight will both suppress neuron $t_{1}$ and also ensure the second leg of the triangle function reaches 00 at $x=1$. This is now the first network shown in Figure 2.
To compose the triangle waves $W_{i}(x)$, copies of this network can be stacked together depth-wise, lining up the “in” and “out” neurons of successive networks. This will have a non-uniform width ($1\rightarrow 2\rightarrow 1\rightarrow 2\rightarrow 1$), but since the input weights of the hidden $t_{1}$ and $t_{2}$ neurons are both $1$, it can be simplified into the form shown on the top right of Figure 2. The weights $1/a$ and $-1/(a-a^{2})$ give the correct way to combine the $t_{1}$ and $t_{2}$ neurons in order to obtain a triangle function, so every hidden neuron will combine the previous $t_{1}$ and $t_{2}$ neuron in this proportion. Using identical weights makes each layer act as if it is composing functions in one dimension, rather than in a dimension equal to the number of neurons.
The goal of the network is not to output triangle waves; instead, the function we seek to approximate is $F(x)$, defined earlier to be a weighted sum of the triangle waves from each layer. We add an extra neuron to each layer, labeled “sum” in Figure 2, to compute this. These are ordinary neurons, but thanks to their specific weights, they will act similarly to a residual connection (He et al., 2016), allowing the hidden features in the network to be used in the output. Each sum neuron adds the previous layer’s triangle wave to the previous sum neuron, thereby iteratively updating an approximation to $F(x)$.
According to Theorem 3, each triangle wave added to the sum must have an amplitude exponentially smaller than the last. Exponentially small weights would pose conditioning issues in optimization, so rather than giving the sum neuron an exponentially small weight, we use the ratio of successive scaling coefficients $S_{i}=s_{i}/s_{i-1}$ to iteratively decay the amplitude of the $t_{1}$ and $t_{2}$ neurons. But because $t_{2}$ now has an exponentially small amplitude, its bias must be exponentially small to function properly. Rather than implementing an exponentially small bias, we make the bias into a neuron. This is rather unconventional, but by using the connection to the previous bias neuron, the bias can decay iteratively, and $t_{2}$ neurons can use their weight to the bias neuron to implement their bias. In matrix form, our hidden layers are:
$\begin{bmatrix}1&\pm[S_{i}/a_{i}&-S_{i}/(a_{i}-a_{i}^{2})]&0\\ 0&S_{i}/a_{i}&-S_{i}/(a_{i}-a_{i}^{2})&0\\ 0&S_{i}/a_{i}&-S_{i}/(a_{i}-a_{i}^{2})&-S_{i}a_{i+1}\\ 0&0&0&S_{i}\end{bmatrix}\times\begin{bmatrix}\text{sum}\\ t_{1}\\ t_{2}\\ \text{bias}\end{bmatrix}.$
Here, rows correspond to neurons, and the sum neuron can either add or subtract the triangle waves depending on whether the target function is convex or concave.
@@ 3.2 Differentiable Model Output
Given the rapid oscillations of the triangle waves formed at each layer, the network will output a fractal with many choices of scaling parameters. This would be poorly predictive of unseen data points generated by a smooth curve. Rather than turning to a general regularizer like batch norm (Ioffe & Szegedy, 2015) or layer norm (Lei Ba et al., 2016), we use a regularization scheme that arises from the form of our functions. When we require that the network output approach a differentiable function in the infinite depth limit, an elegant relationship arises where the peak locations of the triangle waves uniquely determine the scales with which to sum them.
@@
$s_{i+1}=s_{i}(1-a_{i+1})a_{i+2}.$ (3)
Mathematical analysis and sufficient conditions for differentiability in the limit are in Appendix A.3. In Section 4, we will see that following this regularization allows the pretraining phase to produce better initial solutions.
@@ 4 One-Dimensional Experiments
Since our network construction applies directly to one-dimensional convex functions, we focus on such functions in this section (we cover more difficult functions in Sections 56 and Appendices B.2B.3). The aim of these experiments is twofold: (1) we would like to determine the most effective function representations possible from such a small network, and (2) to explore how the utilization of an increased number of linear regions can affect a network’s ability to capture underlying nonlinearity in its training data. To demonstrate that our networks can learn function representations that better utilize depth, we benchmark against PyTorch’s (Paszke et al., 2019) default settings (nn.linear() uses Kaiming initialization), as well as the RAAI distribution (Singh & Sreejith, 2021), and produce errors that are orders of magnitude lower than both. We also train several permutations of our reparameterized networks, which validate that our pretraining and our differentibility constraints indeed facilitate smoother navigation of the loss landscape. Lastly, we conduct a second round of tests to determine if pretrained networks display an enhanced predictive capacity on unseen data points, as might be expected if they can leverage greater nonlinearity in their outputs.
@@ 4.1 Experimental Setup
All models are trained using Adam (Kingma & Ba, 2017) as the optimizer with a learning rate of 0.001 for 1,000 epochs to ensure convergence (see Appendix B.1 for a brief analysis of different learning rates). Each network is four neurons wide with five hidden layers, along with a one-dimensional input and output. The loss function used is the mean squared error, and the average and minimum loss are recorded for 30 models of each type. To compare how well each setup navigates the loss landscape, the networks using triangle-based parameterizations share a common set of starting locations where the triangle peaks and scaling parameters are chosen according to Theorem 3.1. The “unregularized” network has extra degrees of freedom to choose the scaling coefficients after initialization (so that differentiability of the output is not enforced during pretraining). “pretraining skipped” is a network that initializes with our method, but only trains in the standard parameterization. The four curves we approximate are $x^{3}$, $x^{11}$, $\tanh(3x)$, and a quarter period of a sine wave. To approximate the sine and the hyperbolic tangent, the triangle waves are added to the line $y=x$, while for $x^{3}$ and $x^{11}$ the waves are subtracted. When the waves are subtracted, the first scaling factor has to be changed to $a_{0}*a_{1}$ instead of $(1-a_{0})*a_{1}$. The first set of data is 500 evenly spaced points on the interval $[0,1]$ for each of the curves. This is chosen to be very dense deliberately, to try to evoke the most accurate function interpolations from each network. We determine from these tests that pretraining with differentiability enforced produces the best results, so we compare it to standard networks in our second set of experiments. We use a second set of data consisting of only 10 points, with a test set of 10 points spaced in between so as to be as far away from learned data as possible. This set of experiments evaluates the predictive capacities of the networks on unseen data.